PowerDrive, Inc. produces a hard disk drive that sells for $175 per unit. The cost of producing 25,000 drives in the prior year was:
Direct material $625,000
Direct labor 375,000
Variable overhead 125,000
Fixed overhead 1,500,000
Total cost $2,625,000
At the start of the current year, the company received an order for 3,000 drives from a computer company in China. Management of PowerDrive has mixed feelings about the order. On the one hand they welcome the order because they currently have excess capacity. Also, this is the company’s first international order. On the other hand, the company in China is willing to pay only $135 per unit.
What will be the effect on profit of accepting the order?

Round to two decimal places.

Your Answer:
Answer #1

PowerDrive, Inc. produces a hard disk drive that sells for $175 per unit.

The cost of producing 25,000 drives in the prior year was -

Direct material cost = $625,000

Direct labor cost = $375,000

Variable overhead = $125,000

Total variable cost = $625,000 +$375,000 +$125,000 = $1,125,000

Variable cost per unit = Total variable cost/ number of units

=$1,125,000/ 25,000

= $45 per unit

Fixed overhead = $1,500,000

Therefore total cost of producing 25,000 units is $2,625,000

Total Revenue from selling 25,000 units = selling price per unit * number of units

= $175 *25,000 = $4,375,000

And gross profit before accepting international order = Total Revenue from selling 25,000 units - total cost of producing 25,000 units

= $4,375,000 - $2,625,000 = $1,750,000

At the start of the current year, the company received an order for 3,000 drives from a computer company in China.

The company in China is willing to pay only $135 per unit.

This order will increase the variable cost of production but fixed cost will remain same as the company has excess capacity.

Therefore total cost of producing (25,000+3000=) 28,000 units = Total variable cost + total fixed cost

Total variable cost = variable cost per unit * number of units produced

= $45 * 28,000 = $1,260,000

Total Fixed cost/ overhead = $1,500,000

Therefore total cost = $1,260,000 +$1,500,000 = $2,760,000

Total Revenue from selling 28,000 units = selling price per unit for 25,000 units * number of units + selling price per unit for 3,000 units * number of units

= $175 *25,000 + $135 * 3000= $4,375,000 + $405,000

= $4,780,000

And gross profit after accepting international order = Total Revenue from selling 28,000 units - total cost of producing 28,000 units

= $4,780,000 - $2,760,000 = $2,020,000

Therefore the effect on profit of accepting the order = gross profit after accepting international order - gross profit before accepting international order

= $2,020,000 - $1,750,000

= $270,000.00

The profit will increase by $270,000.00